# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

### v1

@@ Abstract
Prompt optimization (PO) offers a practical alternative to fine-tuning large language models (LLMs), enabling performance improvements without altering model weights. Existing methods typically rely on advanced, large-scale LLMs like GPT-4 to generate optimized prompts. However, due to limited downward compatibility, verbose, instruction-heavy prompts from advanced LLMs can overwhelm lightweight inference models and degrade response quality. In this work, we rethink prompt optimization through the lens of interpretable design. We first identify a set of model-agnostic prompt quality merits and empirically validate their effectiveness in enhancing prompt and response quality. We then introduce MePO, a merit-guided, lightweight, and locally deployable prompt optimizer trained on our preference dataset built from merit-aligned prompts generated by a lightweight LLM. Unlike prior work, MePO avoids online optimization reliance, reduces cost and privacy concerns, and—by learning clear, interpretable merits—generalizes effectively to both large-scale and lightweight inference models. Experiments demonstrate that MePO achieves better results across diverse tasks and model types, offering a scalable and robust solution for real-world deployment [^1] Our model and dataset are available at https://github.com/MidiyaZhu/MePO.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing tasks (Achiam et al., 2023; Touvron et al., 2023). However, their output quality remains highly sensitive to prompt phrasing, making effective prompt design crucial for achieving good responses. This has led to growing interest in prompt optimization (PO) (Liu et al., 2023), which aims to improve model performance by refining the input prompt rather than modifying model parameters (Yang et al., ; Zhu et al., 2023).
(a) Prompts from stronger optimization models can degrade the performance of smaller inference models.
(b) Lightweight LLMs fail under complex prompt optimization algorithms.
@@
Figure 1: Empirical analysis of prompt optimization behavior across model scales and optimization algorithms. (a) Raw and GPT-4 optimized prompts are drawn from the Alpaca dataset and Lu et al. (2025). (b) Basic and EvoPrompt examples, along with the optimization algorithm used, are adapted from Guo et al. (2024). (c) The optimized prompt is generated using our merit-guided instruction (Fig. 14). ‘DS’ refers to DeepSeek-R1. Further details are provided in Appx.A.
@@
A dominant PO paradigm is discrete automatic prompt optimization, which employs an advanced, large-scale LLM [^2] In this work, we categorize LLMs like GPT-4 or 70B models as large-scale, while those with fewer than 13B are denoted as lightweight, following ChunLiu et al. (2024). as the prompt optimization model to generate optimized prompts. These prompts are then used as inputs to either similarly sized (Xiang et al., 2025; Ye et al., 2024) or smaller (Guo et al., 2024) inference models for response generation. One line of work adopts online optimization, using advanced API-based models like GPT-4 (Wang et al., ; Zhou et al., 2022; Pryzant et al., 2023). To reduce inference costs, most works typically produce a single optimized task-specific meta-prompt, inserted before or after each query, while leaving the query itself unoptimized (Xiang et al., 2025; Guo et al., 2024; Ye et al., 2024). However, this strategy requires prior knowledge of downstream tasks—an assumption that rarely holds in open-ended, real-world scenarios.
@@
To address this issue, another line of research explores locally deployable prompt optimizers (Lu et al., 2025; Cheng et al., 2024). These models are trained independently of any specific task, though their training data still originates from prompts generated by advanced online LLMs, implicitly treating these outputs as optimal.
@@
Despite their differences in deployment strategies, both online and local optimization approaches rely on prompts produced by powerful online LLMs as the gold-standard optimal prompts. This design implicitly assumes downward compatibility: the notion that prompts generated by advanced LLMs can be effectively interpreted by lightweight inference models. However, our empirical analysis challenges this assumption. As shown in Fig.1(a), prompts crafted by advanced, large-scale LLMs often contain verbose, instruction-heavy content, which may overwhelm lightweight models and degrade response quality.
@@
This finding on the limited downward compatibility of advanced LLMs raises a crucial question: Could lightweight LLMs themselves serve as effective optimization models?
@@
Lightweight models are typically not used for prompt optimization due to their perceived limitations in handling complex optimization processes Wang et al. ; Lu et al. (2025), as illustrated in Fig.1(b). However, our research reveals a promising alternative. We find that these models can in fact serve as effective prompt optimizers when guided by explicit and interpretable prompt design merits (Fig. 1(c)). This suggests that, when equipped with clear, learnable structures, even lightweight LLMs can function as capable optimizers—offering a cost-efficient and locally deployable alternative to online methods, particularly in low-resource or privacy-sensitive settings.
@@
To this end, we propose MePO, a Merit-Guided Prompt Optimization model. We begin by conducting a systematic empirical analysis to identify four sets of interpretable prompt merits that consistently characterize high-quality prompts, offering actionable insights into effective prompt construction. Guided by these merits, we build a prompt optimization preference dataset using optimized prompts generated by a lightweight LLM—eschewing reliance on advanced online LLMs as in prior work—and use it to train an end-to-end prompt optimization model. At the prompt level, our approach produces precise and clear merit guidance that can be generated by lightweight prompt optimization models. At the response level, our optimized prompts not only enhance the performance of similarly scaled inference models, but also demonstrate strong downward and upward compatibility, as the learned prompt merits generalize effectively across both large-scale and lightweight inference models. Besides, constructing the dataset using lightweight models greatly reduces API cost and external dependencies, enabling scalable, privacy-preserving, and locally deployable prompt optimization [^3] While we do not explicitly explore model-adaptive PO in this work, we posit that using the same model architecture for both the optimization and inference models may offer advantages due to shared internal representations..
@@
3. We propose MePO, a lightweight, locally deployable prompt optimization model trained on the merit-aligned dataset.
@@ 2 Empirical Analysis: What Merits a Good Prompt?
Prior APO frameworks commonly rely on advanced online LLMs (Guo et al., 2024; Yuksekgonul et al., 2024; Cheng et al., 2024) to generate optimized prompts. This approach is rooted in the assumption that advanced, large-scale LLMs are inherently better at crafting high-quality prompts (Pan et al., 2023). However, some studies raise concerns about the transferability of such prompts to smaller models. Wang et al. cautions that PromptAgent is designed to optimize prompts for state-of-the-art LLMs, yet these expert-level prompts often fail to transfer to smaller models such as GPT-2 or LLaMA, resulting in substantial performance degradation. Similarly, Lu et al. (2025) observes that small optimizers (13b) fail in difficult prompt optimization tasks.
We argue that these limitations are not solely due to the generation capacity of lightweight LLMs, but rather to the overly vague or underspecified optimization instructions used in prior work. For example, prompts such as “Given error feedback, generate a better prompt” (Wang et al., ), or “You are an expert in prompt optimization” (Lu et al., 2025), often lack concrete guidance. We find that when lightweight LLMs are explicitly instructed with detailed prompt patterns, they can generate effective optimized prompts despite their limited capacity (Fig. 1(c)). This observation motivates a key question: What merits a good prompt?
@@ 2.1 Merit Discovery
Prior studies explore various merits that contribute to prompt effectiveness. Wei et al. (2022) proposes that including Chain-of-Thought (CoT) reasoning in prompts enhances performance, while Lampinen et al. (2022) finds that Detailed Explanations improve prompt quality. Bsharat et al. (2023) introduces 26 prompting principles, categorized into dimensions such as Prompt Structure and Clarity, and Specificity and Information. Ye et al. (2024) proposes three key components for complex reasoning prompts: Detailed Descriptions, Context Specification, and Step-by-Step Reasoning.
@@
Empirical Analysis: We randomly select 1,400 raw questions from the Alpaca dataset Taori et al. (2023) and ask a lightweight LLM [^4] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct to rewrite each question five times [^5] Implementation details are provided in Appx.B.. We then use the same LLM as the inference model to generate responses for both the raw and rewritten questions. The quality of each response is assessed using DeepSeek-R1 (Guo et al., 2025)5. Interestingly, we observe that even small changes in a prompt—sometimes just a few words—can lead to significant differences in response scores. This suggests that certain rewrites yield better prompts, which in turn lead to more effective answers.
Based on this observation, we identify raw-rewritten prompt pairs and categorize them as: (1) prompts that led to higher-scoring responses, and (2) those that resulted in lower-scoring responses, with a score difference greater than 4 points. We then ask DeepSeek-R1 to identify the merits that make the higher-scoring prompts more effective compared to the lower-scoring ones5.
@@
Fig. 2 shows the top 10 merits most frequently associated with high-scoring prompts. Among them, several merits exhibit notably higher frequencies, revealing that high-quality prompts consistently reflect a set of core merits that can be effectively generated by lightweight LLMs. Motivated by this finding, we propose the following design merits to guide the construction of optimal prompts:
@@
$\bullet$ Preserve Original Information: Ensure no information or intent from the original prompt is lost or omitted in the transformation.
Response-Level Evaluation: To assess whether merit-aligned prompts improve response4 quality, we compare responses generated from merit-guided optimized prompts against those from raw prompts across two datasets (2,000 samples each), with scores assigned by DeepSeek-R15.
$\bullet$ Alpaca Dataset [^6] https://huggingface.co/datasets/tatsu-lab/alpaca: Raw responses are generated by text-davinci-003.
$\bullet$ BPO Dataset [^7] https://huggingface.co/datasets/THUDM/BPO: Raw responses are sourced from human-preferred answers.
Figure 3: Response-level win rate comparison. ‘Win’ indicates that the optimized prompt’s response received a higher score than that from the raw prompt.
Figure 4: Alignment Stages of Merit-Guided Prompt Optimization (MePO)
@@
Prompt-Level Evaluation: To move beyond response quality and directly assess prompt quality—a dimension largely overlooked in prior work—we randomly interleaved optimized and raw prompts and asked DeepSeek-R1 to evaluate them in isolation (i.e., without seeing the generated responses)5. Across 2000 prompt comparisons per dataset, optimized prompts were preferred in 95.75% of cases on BPO and 97.9% on Alpaca. These results reinforce the effectiveness of our proposed prompt merits and demonstrate that prompt quality can be reliably evaluated independently of generated outputs.
Conclusion: Our evaluations demonstrate that optimized prompts grounded in our proposed interpretable merits not only yield stronger responses but also exhibit higher prompt quality. Moreover, the results show that effective prompt optimization is not exclusive to advanced, large-scale LLMs such as GPT-4. With merit-based guidance, lightweight LLMs can also serve as capable prompt optimizers, offering scalable and privacy-friendly alternatives for real-world deployment.
@@ 3 Method: From Merits to Optimizer
Our objective is to end-to-end optimize arbitrary input prompts $P_{\text{silver}}$ into refined prompts $P_{\text{golden}}$, enabling inference models to generate higher-quality responses. To achieve this, we introduce MePO, a lightweight LLM-based prompt optimization model designed for local deployment. MePO is trained under a preference learning framework, where supervision is provided by prompt constructed according to the merits identified in Sec. 2. An overview of MePO is shown in Fig. 4.
@@ 3.1 Constructing the Merit-Aligned Prompt Preference Dataset
To train MePO, we construct a Prompt Optimization Preference (POP) dataset grounded in interpretable prompt merits. Following prior work (Cheng et al., 2024; Lu et al., 2025), we adopt Alpaca and BPO as our base datasets. The instruction field in Alpaca and the prompt field in BPO are treated as raw prompts $P_{\text{silver}}$, which are refined using our defined merits to generate $P_{\text{golden}}$.
We use a base model $M_{\text{ref}}$ to generate responses $\hat{R}_{\text{golden}}$ from each optimized prompt $P_{\text{golden}}$, and retrieve the corresponding raw responses $\hat{R}_{\text{silver}}$ from Alpaca’s output and BPO’s bad_res. Each response is then evaluated by DeepSeek-R1, and we retain only those pairs that satisfy two criteria: (1) $\hat{R}_{\text{golden}}$ receives a higher evaluation score than $\hat{R}_{\text{silver}}$, and (2) the score of $\hat{R}_{\text{golden}}$ exceeds 8999According to our scoring rubric in Fig. 13, responses with scores above 8 are considered accurate and well-aligned with the prompt.. The retained samples are relabeled as $R_{\text{golden}}$ and $R_{\text{silver}}$, forming a structured 4-tuple used in training:
@@
Table 1: Overview of our POP dataset. The top section reports prompt statistics paired from Alpaca and BPO, including 10% degraded examples. The bottom section shows validation results from DeepSeek-R1 and GPT-4o, with win rate denoting the proportion where $R_{\text{golden}}$ or $P_{\text{golden}}$ outperforms its silver counterpart.

### v2

@@ Abstract
Prompt optimization (PO) provides a practical way to improve response quality when users lack the time or expertise to manually craft effective prompts. Existing methods typically rely on advanced, large-scale LLMs like GPT-4 to generate optimized prompts. However, due to limited downward compatibility, verbose, and instruction-heavy prompts from advanced LLMs can overwhelm lightweight inference models and degrade response quality. In this work, we rethink prompt optimization through the lens of explicit and interpretable design. We first identify a set of model-agnostic prompt quality merits and empirically validate their effectiveness in enhancing prompt and response quality. We then introduce a merit-guided prompt optimizer (MePO), which is locally deployable and trained on our preference dataset built from merit-guided prompts generated by a lightweight LLM. Unlike prior work, MePO avoids online optimization reliance, reduces cost and privacy concerns, and—by learning clear, interpretable merits—generalizes effectively to both large-scale and lightweight inference models. Experiments demonstrate that MePO achieves better results across diverse tasks and model types, offering a scalable and robust solution for real-world deployment. The code and dataset can be found in https://github.com/MidiyaZhu/MePO.
@@ 1 Introduction
Large language models (LLMs) have achieved impressive results across many NLP tasks (Achiam et al., 2023; Touvron et al., 2023), but their performance remains highly sensitive to prompt phrasing. Although careful prompt crafting can improve output quality, it is often impractical in real-world settings, where users typically lack the time or expertise to manually refine prompts. This has sparked growing interest in automatic prompt optimization (APO) Liu et al. (2023).
(a) Prompts optimized by advanced models can trigger incorrect responses from lightweight inference models. Irrelevant response examples are provided in Appx. A.1.
(b) Lightweight LLMs fail under complex optimization algorithms, yielding results equal to or worse than the Basic Prompt.
@@
Figure 1: Empirical analysis of prompt optimization behavior across model scales and optimization algorithms. (a) Raw and GPT-4 optimized prompts are drawn from the GSM8K Cobbe et al. (2021) and Lu et al. (2025). (b) Basic and EvoPrompt examples, along with the optimization algorithm used, are adapted from Guo et al. (2024). (c) The optimized prompt is generated using our merit-guided instruction (Fig. 14). We present DeepSeek-R1 (DS)’s prompt and response evaluations (Results are consistent with GPT-4o). Further details are provided in Appx.A.
@@
A dominant APO paradigm is discrete automatic prompt optimization, which employs an advanced, large-scale LLM [^1] In this work, we categorize LLMs ranging from GPT-4-size to 70B as large-scale, while those with fewer than 13B are denoted as lightweight, following ChunLiu et al. (2024). as the prompt optimization model to optimize prompts. These prompts are then used as inputs to either similarly sized (Xiang et al., 2025; Ye et al., 2024) or smaller (Guo et al., 2024) inference models for response generation.
@@
One line of work uses API-based models such as GPT-4 for online prompt optimization, typically inserting a task-specific meta-prompt before or after each query to reduce cost, while leaving the query itself unoptimized (Xiang et al., 2025; Guo et al., 2024; Ye et al., 2024). However, this approach requires prior task knowledge, limiting its applicability in open-ended, real-world scenarios. Another line explores locally deployable optimizers that optimize each prompt and are trained independently of specific tasks, though their training data still originates from prompts generated by advanced online LLMs, implicitly treating these outputs as optimal (Lu et al., 2025; Cheng et al., 2024).
@@
Despite the differences in deployment strategies, both online and local optimization approaches rely on prompts produced by advanced online LLMs as the gold-standard optimal prompts. This design implicitly assumes downward compatibility: the notion that prompts generated by large-scale LLMs can be effectively interpreted by lightweight inference models. However, our empirical analysis challenges this assumption. As shown in Fig. 1(a), prompts crafted by advanced, large-scale LLMs often contain verbose, instruction-heavy content, which may overwhelm lightweight models, causing them to miss the key point or engage in excessive reasoning, ultimately leading to incorrect or irrelevant responses.
@@
The finding on the limited downward compatibility of large-scale LLMs raises a crucial question: Can lightweight LLMs themselves serve as effective optimization models?
@@
Lightweight models are typically not used for prompt optimization due to their perceived limitations in handling complex optimization processes Wang et al. (2024); Lu et al. (2025), as further illustrated in Fig. 1(b). However, our research reveals a promising alternative. We find that these models can indeed serve as effective prompt optimizers when guided by explicit and interpretable prompt design merits (Fig. 1(c)). This suggests that, with clear, learnable structures, even lightweight LLMs can function as capable optimizers, offering a cost-efficient and locally deployable alternative to online methods, particularly in low-resource or privacy-sensitive settings.
@@
To this end, we propose MePO, a Merit-Guided Prompt Optimization model. We begin by conducting a systematic empirical analysis to identify four sets of interpretable prompt merits that consistently characterize high-quality prompts, offering actionable insights into effective prompt construction. Guided by these merits, we build a prompt optimization preference dataset using optimized prompts generated by a lightweight LLM—eschewing reliance on advanced online LLMs as in prior work—and use it to train an end-to-end prompt optimization model. At the prompt level, our approach produces precise and clear merit guidance that can be generated by lightweight prompt optimization models. At the response level, our optimized prompts not only enhance the performance of similarly scaled inference models, but also demonstrate strong downward and upward compatibility, as the learned prompt merits generalize effectively across both large-scale and lightweight inference models. Besides, constructing the dataset using lightweight models greatly reduces API cost and external dependencies, enabling scalable, privacy-preserving, and locally deployable prompt optimization.
@@
3. We propose MePO, a lightweight, locally deployable prompt optimization model trained on our merit-guided preference dataset.
@@ 2 Empirical Analysis: What Merits a Good Prompt?
Prior APO frameworks commonly rely on advanced online LLMs (Guo et al., 2024; Yuksekgonul et al., 2024; Cheng et al., 2024) to generate optimized prompts, directly leveraging the high-quality prompt generation capabilities of large-scale models (Pan et al., 2023). However, some studies raise concerns about the transferability of such prompt optimization capabilities to smaller models. Wang et al. (2024) caution that PromptAgent is designed to optimize prompts for state-of-the-art LLMs, yet these expert-level prompts often fail to transfer to smaller models such as GPT-2 or LLaMA2-7b, resulting in substantial performance degradation. Similarly, Lu et al. (2025) observe that small optimizers (13b) fail in difficult prompt optimization tasks.
We argue that these limitations are not solely due to the generation capacity of lightweight LLMs, but rather to the overly vague or underspecified optimization instructions used in prior work. Specifically, Wang et al. (2024) adopt the prompt “Given error feedback, generate a better prompt,” and Lu et al. (2025) use “You are an expert in prompt optimization.” Both are overly general and lack concrete guidance. In contrast, we find that when lightweight LLMs are explicitly instructed with detailed prompt patterns, they can generate effective optimized prompts despite their limited capacity (Fig. 1(c)). This observation motivates a key question: What merits a good prompt?
@@ 2.1 Merit Discovery
Prior studies explore various merits that contribute to prompt effectiveness. Wei et al. (2022) propose that including Chain-of-Thought (CoT) reasoning in prompts enhances performance, while Lampinen et al. (2022) find that Detailed Explanations improve prompt quality. Bsharat et al. (2023) introduce 26 prompting principles, categorized into dimensions such as Prompt Structure and Clarity, and Specificity and Information. Ye et al. (2024) propose three key components for complex reasoning prompts: Detailed Descriptions, Context Specification, and Step-by-Step Reasoning.
@@
Empirical Analysis: We randomly select 5000 raw questions from the Alpaca dataset Taori et al. (2023) and ask a lightweight LLM [^2] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct to rewrite each question five times. We then use the same LLM as the inference model to generate responses for both the raw and rewritten questions. The quality of each response is scored using DeepSeek-R1 (Guo et al., 2025). Interestingly, we observe that even small changes in a prompt—sometimes just a few words—can lead to significant differences in response scores. This suggests that certain rewrites yield better prompts, which in turn lead to more effective answers.
Based on this observation, we identify raw-rewritten prompt pairs and categorize them as: (1) prompts that led to higher-scoring responses, and (2) those that resulted in lower-scoring responses, with a score difference greater than 4 points. We then ask DeepSeek-R1 to identify the merits that make the higher-scoring prompts more effective compared to the lower-scoring ones.
@@
Fig. 2 shows the top 10 merits most frequently associated with high-scoring prompts. Among them, several merits exhibit notably higher frequencies, revealing that high-quality prompts consistently reflect a set of core merits that can be effectively generated by lightweight LLMs. Motivated by this finding and prior insights, we propose the following design merits to guide the construction of optimal prompts:
@@
$\bullet$ Preserve Original Information: Focus on the original prompt, ensuring no information or intent is lost or omitted in the transformation.
Guided by the discovered merits, we optimize the prompts into their merit-guided versions using a lightweight LLM under the instruction in Fig.14.
Response-Level Evaluation: To assess whether merit-guided prompts improve response quality, we compare responses generated from merit-guided prompts against those from raw prompts across two datasets (2,000 samples each), with scores assigned by DeepSeek-R1.
$\bullet$ Alpaca Dataset (30k) [^3] https://huggingface.co/datasets/tatsu-lab/alpaca: Raw responses are generated by text-davinci-003
$\bullet$ BPO Dataset (13.9k) [^4] https://huggingface.co/datasets/THUDM/BPO: Raw responses are sourced from human-preferred answers.
Figure 3: Response-level win rate comparison. ‘Win’ indicates that the merit-guided prompt’s response received a higher score than that from the raw prompt.
Figure 4: Overall stages of merit-guided prompt optimization (MePO).
@@
Prompt-Level Evaluation: To move beyond response quality and directly assess prompt quality—a dimension largely overlooked in prior work—we randomly interleaved optimized and raw prompts and asked DeepSeek-R1 to evaluate them in isolation (i.e., without seeing the generated responses). Across 2000 prompt comparisons per dataset, optimized prompts were preferred in 95.75% of cases on BPO and 97.9% on Alpaca. These results reinforce the effectiveness of our proposed prompt merits and demonstrate that prompt quality can be reliably evaluated independently of generated outputs.
Summary: Our evaluations demonstrate that optimized prompts grounded in our proposed interpretable merits not only yield stronger responses but also exhibit higher prompt quality. Moreover, the results show that effective prompt optimization is not exclusive to advanced, large-scale LLMs such as GPT-4. With merit-based guidance, lightweight LLMs can also serve as capable prompt optimizers, offering scalable and privacy-friendly alternatives for real-world deployment.
The detailed implementation of merit discovery process is provided in Appx.B.
@@ 3 Method: From Merits to Optimizer
Our objective is to end-to-end optimize arbitrary input prompts $P_{\text{silver}}$ into refined prompts $P_{\text{golden}}$, enabling inference models to generate higher-quality responses. To achieve this, we introduce MePO, a lightweight LLM-based prompt optimization model designed for local deployment. MePO is trained under a preference learning framework, supervised by our proposed dataset, where prompts are constructed according to the merits identified in Sec.2. An overview of MePO is shown in Fig.4.
@@ 3.1 Constructing the Merit-Aligned Prompt Preference Dataset
To train MePO, we construct an API-free Prompt Optimization Preference (POP) dataset grounded in our discovered merits. Following Cheng et al. (2024); Lu et al. (2025), we adopt Alpaca and BPO as our base datasets. The instruction field in Alpaca and the prompt field in BPO are treated as raw prompts $P_{\text{silver}}$, which are then refined by a lightweight LLM $M_{\text{ref}}$2 using the defined merits to generate $P_{\text{golden}}$, eliminating dependence on online LLM optimization.
We then use $M_{\text{ref}}$ to generate responses $\hat{R}_{\text{golden}}$ from each optimized prompt $P_{\text{golden}}$, and retrieve the corresponding raw responses $\hat{R}_{\text{silver}}$ from Alpaca’s output and BPO’s bad_res. Each response is then evaluated by DeepSeek-R1, and we retain only those pairs that satisfy two criteria: (1) $\hat{R}_{\text{golden}}$ receives a higher evaluation score than $\hat{R}_{\text{silver}}$, and (2) the score of $\hat{R}_{\text{golden}}$ exceeds 8666According to our scoring rubric in Fig. 13, responses with scores above 8 are considered accurate and well-aligned with the prompt.. The retained samples are relabeled as $R_{\text{golden}}$ and $R_{\text{silver}}$, forming a structured 4-tuple used in training:
@@
Table 1: Overview of our POP dataset. The top section reports the number of prompt–response 4-tuples derived from Alpaca and BPO. The bottom section shows validation results from DeepSeek-R1 and GPT-4o, with win rate denoting the proportion where $R_{\text{golden}}$ or $P_{\text{golden}}$ outperforms its silver counterpart.